if the product of 3 consecutive odd integers is 315,find the sum of the 2 smaller numbers
Answer by CubeyThePenguin(3113)

consecutive odd integers: (x-2), (x), (x+2)

(x-2)(x)(x+2) = 315

315 = 5 * 63 = 5 * 7 * 9

sum of 2 smaller numbers: 5 + 7 = 12